OpenStudy (anonymous):

The half-life of thorium-229 is 7340 years. How long will it take for a sample of this substance to decay to 20% of its original amount? A=Aoe^kt

OpenStudy (anonymous):

set \[\left(\frac{1}{2}\right)^\frac{t}{7340}=.2\] and solve for \(t\)

OpenStudy (anonymous):

change of base formula gives \[\frac{t}{7340}=\frac{\ln(.2)}{\ln(.5)}\] so \[t=\frac{7340\ln(.2)}{\ln(.5)}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you might notice i did not use \[A=A_0e^{kt}\] you can use that too but it takes longer, because first you have to find \(k\) i only used the numbers that were given

OpenStudy (anonymous):

so the answer is?

OpenStudy (anonymous):

you need a calculator at this point, if you want a decimal

OpenStudy (anonymous):

this is what i get http://www.wolframalpha.com/input/?i=7340*ln%28.2%29%2Fln%28.5%29

OpenStudy (anonymous):

that doesnt help

OpenStudy (anonymous):

so i take .2/.5? i got .4

OpenStudy (anonymous):

thanks